Exercise 3.11 Suppose XI , X2. . . . , Xn are iid with mean 0 and variance 1 Show that Σ-M 0. Vnlog n
Answer #1

As mean is given by :

Mean = (Summation i= 1 to n (Xi) )÷n

So (summation i = 1 to n (Xi)) = mean × n

So the LHS of above equation is :

(Mean × n) ÷ (sqrt (n) * log(n)) = (sqrt(n) × mean)/log(n)

As mean = 0, given

= (Sqrt (n) × 0)/ log(n) = 0